OpenStudy (anonymous):

equation of a line passing through (-3,7) and (7,-8)

OpenStudy (anonymous):

y - y1 = (y2-y1)/(x2-x1) * (x-x1) so y-7 = (-8-7)/(7 +3) * (x+3) simplify. Thats all

OpenStudy (anonymous):

So how basically its a fraction type sort? correct?

OpenStudy (anonymous):

If so, how exactly can I do the top? I can do the bottom no problem but idk about the y-7 part

OpenStudy (anonymous):

so, y-7 = (-15/10) * (x+3) 10 (y-7) = -15(x-3) 10y - 70 = -15x + 45 Therefore, 10y + 15x - 115 =0 Now divide by 5 throughout 2y + 3x - 23 = o

OpenStudy (anonymous):

Wowww that makes a lot more sense... I feel dumb haha. Thank you soooo much